# A Comprehensive Survey of Dataset Distillation

### v3

@@ 1 Introduction
Due to the property of extremely high dimensions in the deep learning regime, the data information is hardly disentangled to specific concepts, and thus distilling numerous high-dimensional data into a few points is not a trivial task. Based on the objectives applied to mimic target data, dataset distillation methods can be grouped into meta-learning framework (Sec. 3) and data matching framework (Sec. 4), and these techniques in each framework can be further classified in a much more detailed manner. In the meta-learning framework, the distilled data are considered hyperparameters and optimized in a nested loop fashion according to the distilled-data-trained model’s risk w.r.t. the target data [36, 11]. The data matching framework updates distilled data by imitating the influence of target data on model training from parameter or feature space [24, 27, 31]. Figure 2 presents these different categories of DD algorithms in a tree diagram.
@@
In this paper, we present a comprehensive survey on dataset distillation, and the main objectives of this survey are as follows: (1) present a clear and systematic overview of dataset distillation; (2) review the recent progress and state-of-the-art algorithms and discuss various applications in dataset distillation; (3) give a comprehensive performance comparison w.r.t. different dataset distillation algorithms; and (4) provide a detailed discussion on the limitation and promising directions of dataset distillation to benefit future studies. We note that there is a concurrent survey on dataset distillation [37]. The most notable difference between [37] and our survey is the taxonomy of DD. It plainly classifies DD algorithms into four different categories of meta-model matching, gradient matching, trajectory matching, and distribution matching, while our hierarchical taxonomy first introduces meta-learning and data matching frameworks, which are then followed by a more fine-grained division. Therefore, our taxonomy provides a more systematic framework to track and summarize the development of DD algorithms and better sheds light on future DD studies.
The rest of this paper is organized as follows. We first provide soem background with respect to DD in Section 2. DD methods under the meta-learning framework are described and comprehensively analyzed in Section 3, and a description of the data matching framework follows in Section 4. Section 5 discusses different types of factorized dataset distillation. Next, we report the performance of various distillation algorithms in Section 6. Applications with dataset distillation are shown in Section 8, and finally we discuss challenges and future directions in Section 9.
@@ 2 Background
For the training algorithm alg, $\texttt{alg}(\mathcal{T},\bm{\theta}^{(0)})$ denote the learned parameters returned by empirical risk minimization (ERM) on the dataset $\mathcal{T}$ with the initialized parameter $\bm{\theta}^{(0)}$:
@@ 3 Meta-Learning Framework
By assuming $\mathcal{R}_{\mathcal{D}}\left({\texttt{alg}(\mathcal{S})}\right)\simeq\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$, the target dataset $\mathcal{T}$ is employed as the validation set in terms of the model $\texttt{alg}(\mathcal{S})$}. In consequent, the objective of DD can be converted to minimize $\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$ to improve the generalization performance of $\texttt{alg}(\mathcal{S})$} in terms of $\mathcal{D}$. To this end, dataset distillation falls into the meta-learning area: the hyperparameter $\mathcal{S}$ is updated by the meta (or outer) algorithm, and the base (or inner) algorithm solves the conventional supervised learning problem w.r.t. the synthetic dataset $\mathcal{S}$ [38]; and the dataset distillation task can be formulated to a bilevel optimization problem as follows:
@@
According to the model and optimization methods used in the inner lop, the meta-learning framework of DD can be further classified into two sub-categories of backpropagation through time time (BPTT) approach and kernel ridge regression (KRR) approach.
@@ 3.1 Backpropagation Through Time Approach
Unlike freezing distilled labels, Sucholutsky and Schonlau [16] extended the work of Wang et al. [11] by learning a soft-label in the synthetic dataset $\mathcal{S}$; i.e., the label $y$ in the synthetic dataset is also trainable for better information compression, i.e., $y_{i}=y_{i}-\alpha\nabla_{y_{i}}\mathcal{L}$ for $(\bm{s}_{i},y_{i})\in\mathcal{S}$. Similarly, Bohdal et al. [15] also extended the standard example distillation to label distillation by solely optimizing the labels of synthetic datasets. Moreover, these researchers provided improvements on the efficiency of long inner loop optimization via (1) iteratively updating the model parameters $\bm{\theta}$ and the distilled labels $y$, i.e., one outer step followed by only one inner step for faster convergence; and (2) fixing the feature extractor of neural networks and solely updating the last linear layer with ridge regression to avoid second-order meta-gradient computation. Although the BPTT framework has been shown to underperform other algorithms, Deng and Russakovsky [18] empirically demonstrated that adding momentum term and longer unrolled trajectory ($200$ steps) in the inner loop optimization can considerably enhance the distillation performance; and the inner loop of model training becomes
@@ 3.2 Kernel Ridge Regression Approach
A similar efficient method was also proposed by Zhou et al. [21], which also focused on solving the last-layer in neural networks with KRR. Specifically, the neural network $f_{\bm{\theta}}=g_{\bm{\theta}_{2}}\circ h_{\bm{\theta}_{1}}$ can be decomposed with the feature extractor $h$ and the linear classifier $g$. Then these coworkers fixed the feature extractor $h$ and the linear classifier $g$ possesses a closed-form solution with KRR; and the the distilled data are accordingly optimized with the MSE loss:
@@ 4 Data Matching Framework
where $D(\cdot,\cdot)$ is a distance function, and $\phi(\cdot)$ maps the dataset $\mathcal{S}$ or $\mathcal{T}$ to other informative spaces, such as gradient, parameter, and feature spaces. In practical implementation, the full datasets of $\mathcal{S}$ and $\mathcal{T}$ are often replaced with random sampled batches of $\mathcal{B}_{\mathcal{S}}$ and $\mathcal{B}_{\mathcal{T}}$ for memory saving and faster convergence.
@@ 4.1 Gradient Matching Approach
where $c$ is the class index and $\mathcal{B}_{\mathcal{S}}^{c}$ and $\mathcal{B}_{\mathcal{T}}^{c}$ denotes the batch examples belong to the $c$-th class. However, according to Lee et al. [25], the class-wise gradient matching pays much attention to the class-common features and overlooks the class-discriminative features in the target dataset, and the distilled synthetic dataset $\mathcal{S}$ does not possess enough class-discriminative information, especially when the target dataset is fine-grained; i.e., class-common features are the dominant. Based on this finding, these coworkers proposed an improved objective function of
@@
to better capture constractive signals between different classes through the sum of loss gradients between classes. A similar approach was proposed by Jiang et al. [47], which employed both intraclass and interclass gradient matching by combining Eq. 25 and Eq. 26. Moreover, these researchers measure the difference of gradients by considering the magnitude instead of only considering the angle, i.e., the cosine distance, and the distance function of Eq. 24 is improved to
@@ 4.2 Trajectory Matching Approach
Although trajectory matching demonstrated empirical success, Du et al. [28] argued that $\bm{\theta}^{(k)}$ is generally induced by training on $\mathcal{S}$ for $k$ epochs during natural learning process, while it was directly initialized with $\bm{\theta}_{\mathcal{T}}^{(k)}$ in trajectory matching, which causes the accumulated trajectory error that measures the difference between the parameters from real trajectories of $\mathcal{S}$ and $\mathcal{T}$. Because the accumulated trajectory error is induced by the mismatch between $\bm{\theta}_{\mathcal{T}}^{(k)}$ and real $\bm{\theta}^{(k)}$ from the natural training, these researchers alleviate this error by adding random noise to when initialize $\bm{\theta}^{(k)}$ to improve robustness w.r.t. $\bm{\theta}_{\mathcal{T}}^{(k)}$.
@@ 4.3 Distribution Matching Approach
Although the above parameter-wise matching shows satisfying performance, Zhao and Bilen [31] visualized the distilled data in two dimension and revealed that there is a large distribution discrepancy between the distilled data and the target data. In other words, the distilled dataset cannot comprehensively cover the data distribution in feature space, as shown in Figure 5(b). Based on this discovery, these researchers proposed to match the synthetic and target data from the distribution perspective for dataset distillation. Specifically, they employed the pretrained feature extractor $\psi_{\bm{v}}$ with the parameter $\bm{v}$ to achieve mapping from the input space to the feature space. The synthetic data are optimized according to the objective function
@@
where $c$ denotes the class index. An illustration of distribution matching is also presented Figure 6. As shown in Eq. 32, distribution matching does not rely on the model parameters and drops bilevel optimization for less memory requirement, whereas it empirically underperforms the above gradient and trajectory matching approaches.
@@
Therefore, the total loss in [32] for dataset distillation is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{f}}+\beta\mathcal{L}_{\text{d}}$, where $\beta$ is a positive factor for balancing the feature matching and discrimination loss. Similar to the discrimination loss, Zhao et al. [34] add a classification loss $\mathcal{L}_{CE}$ as regularization to mitigate less classified synthetic data that caused by the first-order moment mean matching:
@@ 4.4 Discussion
The gradient matching approach can be considered as to be short-range parameter matching, while its backpropagation requires second-order gradient computation. Although the trajectory matching approach make ups for this imperfection, the long-range trajectory introduces a recursive computation graph during meta-gradient backpropagation. Different from matching in parameter space, distribution matching approach employs the feature space as the match proxy and also bypasses second-order gradient computation. Although distribution matching has advantages in scalability w.r.t. high-dimensional data, it empirically underperforms trajectory matching, which might be attributed to the mismatch between the comprehensive distribution and decent distillation performance. In detail, distribution matching achieves DD by mimicking features of the target data, so that the distilled data are evenly distributed in the feature space. However, not all features are equally important and distribution matching might waste the budget on imitating less informative features, thereby undermining the distillation performance.
@@ 5 Factorized Dataset Distillation
Code-based DD aims to learn a series of low-dimensional codes for generating highly informative images through a specific generator. In [33], the vectors that are put into the GAN generator are learned to producing informative images. Concretely, these investigators inverse real examples with a GAN generator and collect corresponding latent codes. Then, the latent vectors are further optimized with the distribution matching algorithm. By this, the optimized latent vectors can induce more informative synthetic examples with the pretrained GAN generator. In addition to use the GAN, Kim et al. [26] employed a deterministic multi-formation function $\texttt{multi-form}(\cdot)$ as the decoder to create synthetic data from fewer condensed data, i.e., the synthetic data are generated by $\mathcal{S}=\texttt{multi-form}(\mathcal{C})$. Then the condensed data $\mathcal{C}$ is optimized in an end-to-end fashion by the gradient matching approach. Besides, the Partitioning and Expansion augmentation strategy adopted in [34] can also be regarded as a non-parametric decoder to generate more synthetic data based on the distilled codes. Concretely, while the distilled codes possess the same dimension as real images, Zhao et al. [34] firstly partition each code into several patches and then expand the small patches into the standard dimension during both dataset distillation and network training periods. Recently, Cazenavette et al. [53] argued that this low cross-architecture transferability comes from direct optimization on the pixel space, which makes the distilled data overfit to specific architectures used in the DD process. With this argument, they regularized the distillation by optimizing the intermediate latent codes that are fed into pretrained generative models for synthetic data generation, which largely increases the cross-architecture performance due to the more realistic synthetic data.
@@
where the one-hot label $\bm{y}\in\mathbb{R}^{C\times 1}$, and $\bm{A}_{i}\in\mathbb{R}^{C\times K}$, and $C$ is the number of classes. By plugging different addressing function $\bm{A}_{i}$ into Eq. 37, a total number of $r$ synthetic images can be generated from the bases for each category. During the distillation process, these two elements of $\mathcal{M}$ and $\mathcal{A}$ were optimized with the meta-learning framework. Different from matrix multiplication in [18], Liu et al. [30] employ hallucinator networks as decoders to generate synthetic data from basis. Specifically, a hallucinator network is consist of three parts of a encoder enc, an affine transformation with trainable scale $\sigma$ and shift $\mu$, and a decoder dec, then a basis $\bm{b}$ is put into the hallucinator network to generate corresponding synthetic image $\bm{s}$ as follows:
@@
where the multiplication is element-wise operation. In addition, to enlarge the knowledge divergence encoded by different hallucinator networks for efficient compression, these researchers proposed an adversarial contrastive constraint that maximizes the divergence of features of different generated synthetic data. A similar code-decoder factorization method is also presented in Lee et al. [35], where they adopted an improved distribution matching to optimize the latent codes and decoders. Recently, Zhang et al. [54] employ the powerful generative model as the decoder and optimize it and corresponding latent codes. Specifically, each low-dimensional latent code can generate synthetic images with different classes by feeding the generative model with corresponding conditional class embedding. By this, the number of learnable parameters is not affected by both the real image resolution and the number of classes, which is promising for tackling the large-scale distillation problem.

### v4

@@ 1 Introduction
Due to the property of extremely high dimensions in the deep learning regime, the data information is hardly disentangled to specific concepts, and thus distilling numerous high-dimensional data into a few points is not a trivial task. Based on the objectives applied to mimic target data, dataset distillation methods can be grouped into meta-learning framework (Sec. 3) and data matching framework (Sec. 4), and these techniques in each framework can be further classified in a much more detailed manner. In the meta-learning framework, the distilled data are considered as hyperparameters and optimized in a nested loop fashion according to the distilled-data-trained model’s risk w.r.t. the target data [36, 11]. The data matching framework updates distilled data by imitating the influence of target data on model training from parameter or feature space [24, 27, 31]. Figure 2 presents these different categories of DD algorithms in a tree diagram.
@@
In this paper, we present a comprehensive survey on dataset distillation, and the main objectives of this survey are as follows: (1) present a clear and systematic overview of dataset distillation; (2) review the recent advances in DD algorithms and discuss various applications in dataset distillation; (3) give a comprehensive performance comparison w.r.t. different dataset distillation algorithms; and (4) provide a detailed discussion on the limitation and promising directions of dataset distillation to benefit future studies. We note that there is a concurrent survey on dataset distillation [37]. The most notable difference between [37] and our survey is the taxonomy of DD. It plainly classifies DD algorithms into four different categories of meta-model matching, gradient matching, trajectory matching, and distribution matching, while our hierarchical taxonomy first introduces meta-learning and data matching frameworks, which are then followed by a more fine-grained division. Therefore, our taxonomy provides a more systematic framework to track and summarize the development of DD algorithms and better sheds light on future DD studies.
The rest of this paper is organized as follows. We first provide some background w.r.t. DD in Section 2. DD methods under the meta-learning framework are described and comprehensively analyzed in Section 3, and a description of the data matching framework follows in Section 4. Section 5 discusses different types of factorized dataset distillation. Next, we report the performance of various distillation algorithms in Section 6. Applications with dataset distillation are shown in Section 8, and finally we discuss challenges and future directions in Section 9.
@@ 2 Background
For the training algorithm alg, $\texttt{alg}(\mathcal{T},\bm{\theta}^{(0)})$ denotes the learned parameters returned by empirical risk minimization (ERM) on the dataset $\mathcal{T}$ with the initialized parameter $\bm{\theta}^{(0)}$:
@@ 3 Meta-Learning Framework
By assuming $\mathcal{R}_{\mathcal{D}}\left({\texttt{alg}(\mathcal{S})}\right)\simeq\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$, the target dataset $\mathcal{T}$ is employed as the validation set in terms of the model $\texttt{alg}(\mathcal{S})$}, and in consequence, the objective of DD can be converted to minimize $\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$ to improve the generalization performance of $\texttt{alg}(\mathcal{S})$} in terms of $\mathcal{D}$. To this end, dataset distillation falls into the meta-learning area: the hyperparameter $\mathcal{S}$ is updated by the meta (or outer) algorithm, and the base (or inner) algorithm solves the conventional supervised learning problem w.r.t. the synthetic dataset $\mathcal{S}$ [38]; and the dataset distillation task can be formulated to a bilevel optimization problem as follows:
@@
According to the model and optimization methods used in the inner loop, the meta-learning framework of DD can be further classified into two sub-categories of backpropagation through time time (BPTT) approach and kernel ridge regression (KRR) approach.
@@ 3.1 Backpropagation Through Time Approach
Unlike freezing distilled labels, Sucholutsky and Schonlau [16] extended the work of Wang et al. [11] by learning a soft-label in the synthetic dataset $\mathcal{S}$: i.e., the label $y$ in the synthetic dataset is also trainable for better information compression, i.e., $y_{i}=y_{i}-\alpha\nabla_{y_{i}}\mathcal{L}$ for $(\bm{s}_{i},y_{i})\in\mathcal{S}$. Similarly, Bohdal et al. [15] also extended the standard example distillation to label distillation by solely optimizing the labels of synthetic datasets. Moreover, these researchers provided improvements on the efficiency of long inner loop optimization via (1) iteratively updating the model parameters $\bm{\theta}$ and the distilled labels $y$, i.e., one outer step followed by only one inner step for faster convergence; and (2) fixing the feature extractor of neural networks and solely updating the last linear layer with ridge regression to avoid second-order meta-gradient computation. Although the BPTT framework has been shown to underperform other algorithms, Deng and Russakovsky [18] empirically demonstrated that adding momentum term and longer unrolled trajectory ($200$ steps) in the inner loop optimization can considerably enhance the distillation performance, and the inner loop of model training becomes
@@ 3.2 Kernel Ridge Regression Approach
A similar efficient method was also proposed by Zhou et al. [21], which also focused on solving the last-layer classifier in neural networks with KRR. Specifically, the neural network $f_{\bm{\theta}}=g_{\bm{\theta}_{2}}\circ h_{\bm{\theta}_{1}}$ can be decomposed with the feature extractor $h$ and the linear classifier $g$. Then these coworkers fixed the feature extractor $h$ and the linear classifier $g$ possesses a closed-form solution with KRR, and the distilled data are accordingly optimized with the MSE loss:
@@ 4 Data Matching Framework
where $D(\cdot,\cdot)$ is a distance function, and $\phi(\cdot)$ maps the dataset $\mathcal{S}$ or $\mathcal{T}$ to other informative spaces, such as gradient, parameter, and feature spaces. In practical implementation, the full datasets of $\mathcal{S}$ and $\mathcal{T}$ are often replaced with randomly sampled batches of $\mathcal{B}_{\mathcal{S}}$ and $\mathcal{B}_{\mathcal{T}}$ for memory saving and faster convergence.
@@ 4.1 Gradient Matching Approach
where $c$ is the class index and $\mathcal{B}_{\mathcal{S}}^{c}$ and $\mathcal{B}_{\mathcal{T}}^{c}$ denotes the batch examples belong to the $c$-th class. However, according to [25], the class-wise gradient matching pays much attention to the class-common features and overlooks the class-discriminative features in the target dataset, and the distilled synthetic dataset $\mathcal{S}$ does not possess enough class-discriminative information, especially when the target dataset is fine-grained, i.e., class-common features are dominant. Based on this finding, these coworkers proposed an improved objective function of
@@
to better capture contrastive signals between different classes through the sum of loss gradients between classes. A similar approach was proposed by Jiang et al. [47], which employed both intraclass and interclass gradient matching by combining Eq. 25 and Eq. 26. Moreover, these researchers measure the difference of gradients by considering the magnitude instead of only considering the angle, i.e., the cosine distance, and the distance function of Eq. 24 is improved to
@@ 4.2 Trajectory Matching Approach
Although trajectory matching demonstrated empirical success, Du et al. [28] argued that $\bm{\theta}^{(k)}$ is generally induced by training on $\mathcal{S}$ for $k$ epochs during the natural learning process, while it was directly initialized with $\bm{\theta}_{\mathcal{T}}^{(k)}$ in trajectory matching, which causes the accumulated trajectory error that measures the difference between the parameters from real trajectories of $\mathcal{S}$ and $\mathcal{T}$. Because the accumulated trajectory error is induced by the mismatch between $\bm{\theta}_{\mathcal{T}}^{(k)}$ and real $\bm{\theta}^{(k)}$ from the natural training, these researchers alleviate this error by adding random noise to when initializing $\bm{\theta}^{(k)}$ to improve robustness w.r.t. $\bm{\theta}_{\mathcal{T}}^{(k)}$.
@@ 4.3 Distribution Matching Approach
Although the above parameter-wise matching shows satisfying performance, Zhao and Bilen [31] visualized the distilled data in two-dimensional plane and revealed that there is a large distribution discrepancy between the distilled data and the target data. In other words, the distilled dataset cannot comprehensively cover the data distribution in feature space, as shown in Figure 5(b). Based on this discovery, these researchers proposed to match the synthetic and target data from the distribution perspective for dataset distillation. Specifically, they employed the pretrained feature extractor $\psi_{\bm{v}}$ with the parameter $\bm{v}$ to achieve mapping from the input space to the feature space. The synthetic data are optimized according to the objective function
@@
where $c$ denotes the class index. An illustration of distribution matching is also presented in Figure 6. As shown in Eq. 32, distribution matching does not rely on the model parameters and drops bilevel optimization for less memory requirement, whereas it empirically underperforms the above gradient and trajectory matching approaches.
@@
Therefore, the total loss in [32] for dataset distillation is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{f}}+\beta\mathcal{L}_{\text{d}}$, where $\beta$ is a positive factor for balancing the feature matching and discrimination loss. Similar to the discrimination loss, Zhao et al. [34] add a classification loss $\mathcal{L}_{CE}$ as regularization to mitigate less classified synthetic data caused by the first-order moment mean matching:
@@ 4.4 Discussion
The gradient matching approach can be considered to be short-range parameter matching, while its backpropagation requires second-order gradient computation. Although the trajectory matching approach makes up for this imperfection, the long-range trajectory introduces a recursive computation graph during meta-gradient backpropagation. Different from matching in parameter space, distribution matching approach employs the feature space as the match proxy and also bypasses second-order gradient computation. Although distribution matching has advantages in scalability w.r.t. high-dimensional data, it empirically underperforms trajectory matching, which might be attributed to the mismatch between the comprehensive distribution and decent distillation performance. In detail, distribution matching achieves DD by mimicking features of the target data, so that the distilled data are evenly distributed in the feature space. However, not all features are equally important and distribution matching might waste the budget on imitating less informative features, thereby undermining the distillation performance.
@@ 5 Factorized Dataset Distillation
Code-based DD aims to learn a series of low-dimensional codes for generating highly informative images through a specific generator. In [33], the vectors that are put into the GAN generator are learned to produce informative images. Concretely, these investigators inverse real examples with a GAN generator and collect corresponding latent codes. Then, the latent vectors are further optimized with the distribution matching algorithm. By this, the optimized latent vectors can induce more informative synthetic examples with the pretrained GAN generator. In addition to using the GAN, Kim et al. [26] employed a deterministic multi-formation function $\texttt{multi-form}(\cdot)$ as the decoder to create synthetic data from fewer condensed data, i.e., the synthetic data are generated by $\mathcal{S}=\texttt{multi-form}(\mathcal{C})$. Then the condensed data $\mathcal{C}$ is optimized in an end-to-end fashion by the gradient matching approach. Besides, the Partitioning and Expansion augmentation strategy adopted in [34] can also be regarded as a non-parametric decoder to generate more synthetic data based on the distilled codes. Concretely, while the distilled codes possess the same dimension as real images, Zhao et al. [34] firstly partition each code into several patches and then expand the small patches into the standard dimension during both dataset distillation and network training periods. Recently, Cazenavette et al. [53] argued that this low cross-architecture transferability comes from direct optimization on the pixel space, which makes the distilled data overfit to specific architectures used in the DD process. With this argument, they regularized the distillation by optimizing the intermediate latent codes that are fed into pretrained generative models for synthetic data generation, which largely increases the cross-architecture performance due to the more realistic synthetic data.
@@
where the one-hot label $\bm{y}\in\mathbb{R}^{C\times 1}$, and $\bm{A}_{i}\in\mathbb{R}^{C\times K}$, and $C$ is the number of classes. By plugging different addressing functions $\bm{A}_{i}$ into Eq. 37, a total number of $r$ synthetic images can be generated from the bases for each category. During the distillation process, these two elements of $\mathcal{M}$ and $\mathcal{A}$ were optimized with the meta-learning framework. Different from matrix multiplication in [18], Liu et al. [30] employ hallucinator networks as decoders to generate synthetic data from basis. Specifically, a hallucinator network consists of three parts of an encoder enc, an affine transformation with trainable scale $\sigma$ and shift $\mu$, and a decoder dec, then a basis $\bm{b}$ is put into the hallucinator network to generate corresponding synthetic image $\bm{s}$ as follows:
@@
where the multiplication is element-wise. In addition, to enlarge the knowledge divergence encoded by different hallucinator networks for efficient compression, these researchers proposed an adversarial contrastive constraint that maximizes the divergence of features of different generated synthetic data. A similar code-decoder factorization method is also presented in Lee et al. [35], where they adopted an improved distribution matching to optimize the latent codes and decoders. Recently, Zhang et al. [54] employ the powerful generative model as the decoder and optimize it and corresponding latent codes. Specifically, each low-dimensional latent code can generate synthetic images with different classes by feeding the generative model with corresponding conditional class embedding. By this, the number of learnable parameters is not affected by both the real image resolution and the number of classes, which is promising for tackling the large-scale distillation problem.